Your firm is considering overhauling its production plant. The engineering team has come up with two proposals, one for a minor overhaul and one for a major overhaul. The two options are mutually exclusive and have the following cash flows. What is the crossover rate? If the cost of capital for both of these projects is 15%, what should your firm do?

Year Cash Flow ($ millions)
Proposal 0 1 2 3
Minor Overhaul -10 6 6 6
Major Overhaul -50 25 25 25
Answer #1

E21 fx А B с D E F 1 2 Years Difference in Cash flows Minor Overhaul Major Overhaul -10 3 0 -50 -401 4 1 6 25 19 5 2 6 25 19E21 for A B с D E 1 Years Minor Overhaul N 3 0 -10 Major Overhaul -50 25 4 1 6 Difference in Cash flows =D3-C3 =D4-C4 =D5-C5